the sides of a triangle measure 6, 3, and 4. If the smallest side of a similar triangle measures 5, what is the largest side?

Answer by TimothyLamb(4379)

x/6 = 5/3
x = 6*5/3
x = 10
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---